OpenStudy (anonymous):

I need to restrict the domain so that f(x)=(x+6)^2 becomes a one-to-one function?

OpenStudy (anonymous):

sorry I never learned this... should I factor or is the answer simply -6?

OpenStudy (solomonzelman):

Do you know what a one-to-one function is?

OpenStudy (anonymous):

yes- there is solely one output for each input, and only one input for one output. passing both the vertical and horizontal line test.

OpenStudy (solomonzelman):

Yes, very nice!

OpenStudy (anonymous):

Sorry I didn't learn how to do set restrictions on the domain to change this to a one-to-one... the exact instruction tells me "restrict the domain of the function so that it is a one-to-one function and had has an inverse

OpenStudy (solomonzelman):

\( \color{#000000 }{ \displaystyle f(x)=(x-h)^2 }\) is symmetrical about a vertical line \(\small \color{#000000 }{ \displaystyle x=h }\). (Right?) Alternatively, it "mirrors" or "reflects" itself about the line \(\small \color{#000000 }{ \displaystyle x=h }\). Of course, this is a problem, because that means that the function fails the horizontal line test.

OpenStudy (solomonzelman):

By, "restricting the domain", you are just setting the domain for your function, where f(x) is a one-to-one function. (i.e. without that "mirror"/"repetition" on the other side)

OpenStudy (anonymous):

would restricting the domain include changing the format of the equation, or would the function remain the same in this instance? can (x+6)^2 go to (x+6)(x-6) for example?

OpenStudy (solomonzelman):